OpenStudy (anonymous):

find the circumfernce. radius=3.5yds

OpenStudy (anonymous):

\[\pi r\]

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

\[C=2 \pi r\]

OpenStudy (anonymous):

C=Circumference r=radius

OpenStudy (anonymous):

oh okay....:)

OpenStudy (anonymous):

Yes. \[2 \pi r\] My bad, I meant \[\pi d\] I apologize.

OpenStudy (anonymous):

its k